OpenStudy (quickstudent):

How can I use FOIL to solve this: (-16t + 1) (x + 8) ?

OpenStudy (kainui):

What does FOIL stand for?

OpenStudy (quickstudent):

first, outside, inside, last.

OpenStudy (mokeira):

First, Outer , inner,Last

OpenStudy (mokeira):

do you know BODMAS?

OpenStudy (quickstudent):

No, I don't know that.

OpenStudy (texaschic101):

FOIL -- first, outside, inside, last (-16t + 1) ( x + 8) First -- multiply first terms in each set....-16t * x = -16tx Outside -- multiply outside terms in each set...-16t * 8 = -128t Inside -- multiply inside terms in each set... 1 * x = x Last -- multiply last terms in each set...1 * 8 = 8 now combine like terms -16tx - 128t + x + 8 <== there you go, your answer

OpenStudy (kainui):

So if you follow what FOIL stands for, just look at these two: (-16t + 1) (x + 8) Which are the First terms? Which are the Outer terms? Inner? Last? Multiply each of these 4 pairs together, and then add them.

OpenStudy (mokeira):

first recognize the first, outer inner and last terms

OpenStudy (quickstudent):

Ok, I understand now.